myininaya (myininaya):

Solve for x: 6x-1 = 9 mod 8 first one to answer correctly is a winner :)

OpenStudy (mimi_x3):

x = 10/6 = 5/3 what's mod 8 ? lols

OpenStudy (mimi_x3):

my bad x = 4n + 3

OpenStudy (anonymous):

6x-1=1 6x=2 x=1/3

myininaya (myininaya):

gj mimi :)

OpenStudy (anonymous):

Mod[9,8]=1

myininaya (myininaya):

= is suppose to be that congruency symbol

OpenStudy (anonymous):

what is mod 0.o?

OpenStudy (anonymous):

this isnt a bad problem either

myininaya (myininaya):

yours is better

myininaya (myininaya):

in number theory we never got to these

myininaya (myininaya):

we only did simple looking stuff lol

OpenStudy (anonymous):

you have taken number theory?

myininaya (myininaya):

you haven't?

myininaya (myininaya):

i loved number theory i wished they offered another number theory class for me to take when i was being schooled

OpenStudy (anonymous):

i have, but i remeber you saying that you were going to take it

myininaya (myininaya):

like a more advanced number theory

myininaya (myininaya):

no not me

myininaya (myininaya):

joe is

OpenStudy (anonymous):

right

myininaya (myininaya):

and someone else

myininaya (myininaya):

i am going to learn it by myself

OpenStudy (anonymous):

zark right?

myininaya (myininaya):

no zarkon has a phd hes done